# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ 5 Conclusion
In this paper, we introduced a novel approach for graph data augmentation that enhances both the generalization and robustness of GNNs. Our method uses Gaussian Mixture Models (GMMs) applied at the output level of the Readout function, an approach motivated by theoretical findings. Using the universal approximation property of GMMs, we can sample new graph representations to effectively control the upper bound of Rademacher Complexity, ensuring improved generalization of Graph Neural Networks (GNNs), as shown in Theorem 3.1. Through extensive experiments on widely used datasets, we demonstrated that our approach not only exhibits strong generalization ability but also maintains robustness against structural perturbations. An additional advantage of our method is its efficiency in terms of time complexity. Unlike baselines that generate augmented data for each individual or pair of training graphs, our approach fits the GMM to the entire training dataset at once, allowing for fast graph data augmentation without incurring significant additional backpropagation time.
@@
In this section, we provide a detailed proof of Theorem 3.1, aiming to derive a theoretical upper bound for both the generalization gap and the Rademacher complexity. Theorem 3.1 Let $\ell$ be a classification loss function with $L_{Lip}$ as a Lipschitz constant such as $\ell(\cdot,\cdot)\in[0,1].$ Then, with a probability at least $1-\delta$ over the samples $\mathcal{D}_{train}$, we have,
$\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]\leq 2\mathcal{R}(\ell_{aug})+5\sqrt{\frac{2log(4/\delta)}{N}}+2L_{Lip}\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right].$
@@
$\mathcal{R}(\ell_{aug})\leq\mathcal{R}(\ell)+\max_{n\in\{1,\ldots,N\}}L_{Lip}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}_{n}^{\lambda}-\mathcal{G}_{n}\right\|\right].$
@@
We will decompose $\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]$ into a finite sum of 5 terms as follows,
$\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]=u_{1}+u_{2}+u_{3}+u_{4}+u_{5}$
@@
$u_{1}$ $=\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\hat{\theta}_{aug})\right]\right],$
$u_{2}$ $=\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\hat{\theta}_{aug})\right]\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\hat{\theta}_{aug})\right],$
$u_{3}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\hat{\theta}_{aug})\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta_{\star})\right],$
$u_{4}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta_{\star})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta_{\star})\right]\right],$
$u_{5}$ $=\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta_{\star})\right]\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right].$
@@
$u_{1}+u_{5}$ $=\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\hat{\theta}_{aug})\right]\right]+\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta_{\star})\right]\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]$
$\leq\left|\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\hat{\theta}_{aug})\right]\right]\right|+\left|\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta_{\star})\right]\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]\right|$
$\leq 2\sup_{\theta\in\Theta}\left|\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta)\right]\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta)\right]\right|$
$\leq 2\sup_{\theta\in\Theta}\left|\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta)\right]-\ell(\mathcal{G},\theta)\right]\right|$
$\leq 2\sup_{\theta\in\Theta}\left|\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta)-\ell(\mathcal{G},\theta)\right]\right]\right|$
$\leq 2L_{Lip}\sup_{\theta\in\Theta}\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|\right].$
For the term $u_{4}$, we apply McDiarmid’s inequality. Since the classification loss satisfy $\ell(\cdot)\in[0,1]$, we get for $k\in\{0,\ldots,N\}$,
$\forall\{(\mathcal{G}_{n},y_{n})\}_{n=1}^{N},\{(\mathcal{G^{\prime}}_{n},y^{\prime}_{n})\}_{n=1}^{N},\theta,$ such that $\forall n\neq k,\quad\mathcal{G}_{n}=\mathcal{G^{\prime}}_{n}$ and $\mathcal{G}_{k}\neq\mathcal{G^{\prime}}_{k}$:
$\left|\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n},\theta)\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$ $=\frac{1}{N}\left|\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$
$\leq\frac{1}{N}\left|\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{k},\theta)-\ell(\mathcal{G}_{k}^{\prime},\theta)\right]\right|$
@@
The first equality is obtained by your claim that $\forall n\neq k,\quad\mathcal{G}_{n}=\mathcal{G^{\prime}}_{n}$ and $\mathcal{G}_{k}\neq\mathcal{G^{\prime}}_{k}$, the last inequality is obtained by the fact that $\ell(\cdot)\in[0,1]$.
@@
$\forall t>0,\quad\mathbb{P}\left(u_{4}\geq t\right)$ $=\mathbb{P}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta_{\star})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\theta_{\star})\right]\right]\geq t\right)$
$\leq exp\left(-\frac{2t^{2}}{\sum_{n=1}^{N}4/N^{2}}\right)$
$=exp\left(-\frac{Nt^{2}}{2}\right).$
Therefore, for $\delta\in]0,1]$, and for $t=\sqrt{2log(1/\delta)/(N)}$, i.e. $exp\left(-\frac{Nt^{2}}{2}\right)=\delta.$, we have,
$\mathbb{P}\left(u_{4}\geq\sqrt{2log(1/\delta)/(N)}\right)\leq\delta.$
@@
$\mathbb{P}\left(u_{4}<\sqrt{\frac{2log(1/\delta)}{N}}\right)=1-\mathbb{P}\left(u_{4}\geq\sqrt{\frac{2log(1/\delta)}{N}}\right)\geq 1-\delta.$
@@
$u_{4}\leq\sqrt{\frac{2log(1/\delta)}{N}}<\sqrt{\frac{2log(4/\delta)}{N}}.$
@@
$u_{2}=\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\hat{\theta}_{aug})\right]\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\hat{\theta}_{aug})\right]\leq 2\mathcal{R}(\ell_{aug})+4\sqrt{\frac{2log(4/\delta)}{N}}.$
The above inequality tells us that the true risk $\mathbb{E}_{\mathcal{G}\sim G}\left[\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}^{\lambda},\hat{\theta}_{aug})\right]\right]$ is bounded by the empirical risk $\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\hat{\theta}_{aug})\right]$ plus a term depending on the Rademacher complexity of the augmented hypothesis class and an additional term that decreases with the size of the sample $N$.
Additionally, since $\hat{\theta}_{aug}$ is the optimal parameter for the loss $\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta)\right]$, thus,
$u_{3}\leq 0$
@@
$\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\hat{\theta}_{aug})\right]-\mathbb{E}_{\mathcal{G}\sim G}\left[\ell(\mathcal{G},\theta_{\star})\right]<2\mathcal{R}(\ell_{aug})+5\sqrt{\frac{2log(4/\delta)}{N}}+2L_{Lip}\mathbb{E}_{\mathcal{G}\sim G}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}_{n}^{\lambda}-\mathcal{G}_{n}\right\|\right].$

### v3

@@ 5 Conclusion
We introduced GRATIN, a novel approach for graph data augmentation that enhances both the generalization and robustness of GNNs. Our method uses Gaussian Mixture Models (GMMs) applied at the output level of the Readout function, an approach motivated by theoretical findings. Using the universal approximation property of GMMs, we can sample new graph representations to effectively control the upper bound of the Rademacher complexity, ensuring improved generalization of GNNs. Through extensive experiments on widely used datasets, we demonstrated that our approach not only exhibits strong generalization ability but also maintains robustness against structural perturbations. An additional advantage of our approach is its efficiency in terms of time complexity. Unlike baselines that generate augmented data for each individual or pair of training graphs, GRATIN fits the GMM to the entire training dataset at once, allowing for fast graph data augmentation without incurring significant additional backpropagation time.
Acknowledgment
Y.A. and M.V. are supported by the French National Research Agency (ANR) via the AML-HELAS (ANR-19-CHIA-0020) project. Y.A. and J.L. are supported by the French National Research Agency (ANR) via the “GraspGNNs” JCJC grant (ANR-24-CE23-3888). F.M. acknowledges the support of the Innov4-ePiK project managed by the French National Research Agency under the 4th PIA, integrated into France2030 (ANR-23-RHUS-0002). This work was granted access to the HPC resources of IDRIS under the allocation “2024-AD010613410R2” made by GENCI.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgment
In this section, we provide a detailed proof of Theorem 3.1, aiming to derive a theoretical upper bound for both the generalization gap and the Rademacher complexity.
Theorem 3.1 Let $\ell$ be a classification loss function with $\text{L}_{\text{Lip}}$ as a Lipschitz constant and $\ell(\cdot,\cdot)\in[0,1].$ Then, with a probability at least $1-\delta$ over the samples $\mathcal{D}_{\text{train}}$, we have,
$\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]\leq 2\mathcal{R}(\ell_{\text{aug}})+5\sqrt{\frac{2\log(4/\delta)}{N}}+2\text{L}_{\text{Lip}}\mathbb{E}_{\mathcal{G}\sim\mathcal{D},\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\left\|\widetilde{\mathcal{G}}-\mathcal{G}\right\|\right].$
@@
$\mathcal{R}(\ell_{\text{aug}})\leq\mathcal{R}(\ell)+\max_{n}\text{L}_{\text{Lip}}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\left\|\widetilde{\mathcal{G}}_{n}^{m}-\mathcal{G}_{n}\right\|\right].$
@@
We will decompose $\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]$ into a finite sum of 5 terms as follows,
$\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]=u_{1}+u_{2}+u_{3}+u_{4}+u_{5}$
@@
$u_{1}$ $=\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\hat{\theta}_{\text{aug}})\right]\right],$
$u_{2}$ $=\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\hat{\theta}_{\text{aug}})\right]\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}_{\text{aug}})\right],$
$u_{3}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}_{\text{aug}})\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}_{\star})\right],$
$u_{4}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}_{\star})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta_{\star})\right]\right],$
$u_{5}$ $=\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta_{\star})\right]\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right].$
@@
$u_{1}+u_{5}$ $=\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\hat{\theta}_{\text{aug}})\right]\right]+\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta_{\star})\right]\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]$
$\leq\left|\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\hat{\theta}_{\text{aug}})\right]\right]\right|+\left|\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta_{\star})\right]\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]\right|$
$\leq 2\sup_{\theta\in\Theta}\left|\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta)\right]\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta)\right]\right|$
$\leq 2\sup_{\theta\in\Theta}\left|\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta)\right]-\ell(\mathcal{G},\theta)\right]\right|$
$\leq 2\sup_{\theta\in\Theta}\left|\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta)-\ell(\mathcal{G},\theta)\right]\right]\right|$
$\leq 2\text{L}_{\text{Lip}}\sup_{\theta\in\Theta}\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\left\|\widetilde{\mathcal{G}}-\mathcal{G}\right\|\right].$
For the term $u_{4}$, we apply McDiarmid’s inequality; we consider the two sets $\{(\mathcal{G}_{n},y_{n})\}_{n=1}^{N}$ and $\{(\mathcal{G}_{n}^{\prime},y_{n}^{\prime})\}_{n=1}^{N}$ are identical except at the $k$-th element, i.e. for a fixed $k\in\{1,\ldots,N\}$ we have $\mathcal{G}_{k}^{\prime}\neq\mathcal{G}_{k}$ and $\forall n\in\{1,\ldots,N\},~{}n\neq k\Rightarrow\mathcal{G}_{n}^{\prime}=\mathcal{G}_{n}$. This setup allows us to bound the change in the expected loss when a single graph is replaced. Since the classification loss satisfy $\ell(\cdot)\in[0,1]$, we get,
$\left|\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\mathcal{G}_{n},\theta)\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$ $=\frac{1}{N}\left|\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$
$\leq\frac{1}{N}\left|\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\mathcal{G}_{k},\theta)-\ell(\mathcal{G}_{k}^{\prime},\theta)\right]\right|$
@@
The first equality is obtained by the fact that $\forall n\neq k,\quad\mathcal{G}_{n}=\mathcal{G^{\prime}}_{n}$ and $\mathcal{G}_{k}\neq\mathcal{G^{\prime}}_{k}$, the last inequality is obtained by the fact that $\ell(\cdot)\in[0,1]$.
@@
$\forall t>0,\quad\mathbb{P}\left(u_{4}\geq t\right)$ $=\mathbb{P}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta_{\star})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\theta_{\star})\right]\right]\geq t\right)$
$\leq\exp\left(-\frac{2t^{2}}{\sum_{n=1}^{N}4/N^{2}}\right)$
$=\exp\left(-\frac{Nt^{2}}{2}\right).$
Therefore, for $\delta\in]0,1]$, and for $t=\sqrt{2\log(1/\delta)/N}$, i.e. $exp\left(-\frac{Nt^{2}}{2}\right)=\delta.$, we have,
$\mathbb{P}\left(u_{4}\geq\sqrt{2\log(1/\delta)/N}\right)\leq\delta.$
@@
$\mathbb{P}\left(u_{4}<\sqrt{\frac{2\log(1/\delta)}{N}}\right)=1-\mathbb{P}\left(u_{4}\geq\sqrt{\frac{2\log(1/\delta)}{N}}\right)\geq 1-\delta.$
@@
$u_{4}\leq\sqrt{\frac{2\log(1/\delta)}{N}}<\sqrt{\frac{2\log(4/\delta)}{N}}.$
@@
$u_{2}=\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\hat{\theta}_{\text{aug}})\right]\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}_{\text{aug}})\right]\leq 2\mathcal{R}(\ell_{\text{aug}})+4\sqrt{\frac{2\log(4/\delta)}{N}}.$
The above inequality tells us that the true risk $\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\mathbb{E}_{\widetilde{\mathcal{G}}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}},\hat{\theta}_{\text{aug}})\right]\right]$ is bounded by the empirical risk $\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}_{\text{aug}})\right]$ plus a term depending on the Rademacher complexity of the augmented hypothesis class and an additional term that decreases with the size of the sample $N$.
Additionally, since $\hat{\theta}_{\text{aug}}$ is the optimal parameter for the loss $\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta})\right]$, thus,
$u_{3}\leq 0.$
@@
$\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\hat{\theta}_{\text{aug}})\right]-\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\left[\ell(\mathcal{G},\theta_{\star})\right]<2\mathcal{R}(\ell_{\text{aug}})+5\sqrt{\frac{2\log(4/\delta)}{N}}+2\text{L}_{\text{Lip}}\mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\mathbb{E}_{\widetilde{\mathcal{G}}_{n}^{m}\sim A_{\lambda}}\left[\left\|\widetilde{\mathcal{G}}_{n}^{m}-\mathcal{G}_{n}\right\|\right].$